Arrival times given the count as uniform order statistics

[22/64%]
Show
LaTeX

Event AA occurs mm times during a time period T0T_{0}, with the probability of its occurrence being the same at every moment. Determine the probability that it occurred kk times during a given interval TT, letting T0→∞T_{0} \to \infty and mT0→p\frac{m}{T_{0}} \to p.

?

This exercise asks you to prove that the Poisson process conditioned to be equal to nn at time 1 has the same law as nn times the empirical process. Here is the precise statement. Suppose PtP_{t} is a Poisson process with parameter λ>0\lambda >0. Let Q\mathbb {Q} be the law of {Pt,t∈[0,1]}\left\{ P_{t}, t \in [0,1]\right\} conditioned so that P1=nP_{1}=n. Thus Q\mathbb {Q} is a probability on D[0,1]D[0,1] with

Q(P∈A)=P(P∈A∣P1=n). \mathbb {Q}(P \in A)=\mathbb {P}\left( P \in A \mid P_{1}=n \right) .

Since (P1=n)\left(P_{1}=n\right) is an event with positive probability, there is no difficulty defining these conditional probabilities. Prove that Q\mathbb {Q} is also the law of {nFn(t),t∈[0,1]}\left\{ n F_{n}(t), t \in [0,1]\right\}, where FnF_{n} is defined in Section 35.3.

?

For the Poisson process (see Example 33.1) show that for 0<s<t0<s<t,

P[Ns=k∥Nt]={(Ntk)(st)k(1−st)Nt−k,k≤Nt0,k>Nt P\left[N_{s}=k \| N_{t}\right]= \begin{cases} \binom {N_{t}}{k}\left(\frac{s}{t}\right)^{k}\left(1-\frac{s}{t}\right)^{N_{t}-k}, & k \leq N_{t} \\ 0, & k>N_{t} \end{cases}

Thus the conditional distribution (in the sense of Theorem 33.3) of NsN_{s} given NtN_{t} is binomial with parameters NtN_{t} and s/ts / t.

?

33.734.11↑33.734 .11 \uparrow Use Example 33.10 to calculate P[Ns=k∥Nu,u≥t](s≤t)P\left[N_{s}=k \| N_{u}, u \geq t\right](s \leq t) for the Poisson process.

?

Let XX follow a Poisson distribution with parameter λ\lambda. We select NN random points from the interval [0,1][0, 1]. Let XiX_{i} be the number of points in the interval [(i−1)/n,i/n][(i-1)/n, i/n] (i=1,…,ni = 1, \ldots , n). Show that the random variables XiX_{i} are independent and identically distributed.

?
?
(a)

Show that the arrival epochs of a Poisson process satisfy

fS→(n)∣Sn+1(s→(n)∣sn+1)=n!/sn+1n. f_{\overrightarrow {S}^{(n)} \mid S_{n+1}}\left(\overrightarrow {s}^{(n)} \mid s_{n+1}\right) = n!/s_{n+1}^n.
(b)

Contrast this with the result of Theorem 2.5.1.

Remark.
?
(a)

This is easy if you use only the results of Section 2.2.2.

Equation (2.42) gives fSi∣N(t)(si∣n)f_{S_i \mid N(t)}(s_i \mid n), which is the density of rv SiS_i conditional on N(t)=nN(t) = n for n≥in \geq i. Multiply this expression by P(N(t)=n)\mathbb {P}\left(N(t) = n\right) and sum over nn to find fSi(si)f_{S_i}(s_i); verify that your answer is indeed the Erlang density.

?
?
(a)

For a Poisson counting process of rate λ\lambda, find the joint probability density of S1,S2,…,Sn−1S_1, S_2, \ldots , S_{n-1} conditional on Sn=tS_n = t.

(b)

Find P(X1>τ∣Sn=t)\mathbb {P}\left(X_1 > \tau \mid S_n=t\right).

(c)

Find P(Xi>τ∣Sn=t)\mathbb {P}\left(X_i > \tau \mid S_n=t\right) for 1≤i≤n1 \leq i \leq n.

(d)

Find the density fSi∣Sn(si∣t)f_{S_i \mid S_n}(s_i \mid t) for 1≤i≤n−11 \leq i \leq n-1.

(e)

Give an explanation for the striking similarity between the condition N(t)=n−1N(t) = n-1 and the condition Sn=tS_n = t.

?
(a)

For a Poisson process of rate λ\lambda, find P(N(t)=n∣S1=τ)\mathbb {P}\left(N(t)=n \mid S_1=\tau \right) for t>τt > \tau and n≥1n \geq 1.

(b)

Using this, find fS1(τ∣N(t)=n)f_{S_1}(\tau \mid N(t)=n).

(c)

Check your answer against (2.41).

?
(a)

Use Equation (2.42) -- fSi∣N(t)(si∣n)=sii−1(i−1)!(t−si)n−i(n−i)!n!tnf_{S_i \mid N(t)}(s_i \mid n) = \frac{s_i^{i-1}}{(i-1)!} \frac{(t-s_i)^{n-i}}{(n-i)!} \frac{n!}{t^n} for n≥in \geq i -- to find E[Si∣N(t)=n]\mathbb {E}\left[S_i \mid N(t)=n\right].

(b)

Find the second moment and the variance of SiS_i conditional on N(t)=nN(t)=n.

(c)

Assume that nn is odd, and consider i=(n+1)/2i = (n+1)/2. What is the relationship between SiS_i, conditional on N(t)=nN(t)=n, and the sample median of nn IID uniform rv's.

(d)

Give a weak law of large numbers (WLLN) for the above median.

Remark.
?
(a)

When you integrate sifSi(si∣N(t)=n)s_i f_{S_i}(s_i \mid N(t)=n), compare this integral with fSi+1(si∣N(t)=n+1)f_{S_{i+1}}(s_i \mid N(t)=n+1) and use the fact that the latter expression is a probability density.

(b)

Extend the previous hint.

?
(a)

For 1≤i<n1 \leq i < n, find the conditional density of Si+1S_{i+1}, conditional on N(t)=nN(t) = n and Si=siS_i = s_i.

(b)

Use (a) to find the joint density of S1,…,SnS_1, \ldots , S_n conditional on N(t)=nN(t) = n. Verify that your answer agrees with (2.38).

Let Sm=∑r=1mXr,m≤nS_{m} = \sum_{r = 1}^{m} X_{r}, m \leq n, where the XrX_{r} are independent and identically distributed with finite mean. Denote by U1,U2,…,UnU_{1}, U_{2}, \ldots , U_{n} the order statistics of nn independent variables which are uniformly distributed on (0,t)(0, t), and set Un+1=tU_{n+1} = t. Show that Rm=Sm/Um+1,0≤m≤nR_{m} = S_{m} / U_{m+1}, 0 \leq m \leq n, is a backward martingale with respect to a suitable sequence of σ\sigma-fields, and deduce that

P(Rm≥1 for some m≤n∣Sn=y)≤min⁡{y/t,1} \mathbb {P}\left(R_{m} \geq 1 \text{ for some } m \leq n \mid S_{n} = y\right) \leq \min \left\{ y / t, 1\right\}
?

Let {Xr:r≥1}\left\{ X_{r}: r \geq 1\right\} be independent exponential random variables with parameter λ\lambda, and set Sn=∑r=1nXrS_{n} = \sum_{r = 1}^{n} X_{r}. Show that:

?
(a)

Yk=Sk/Sn,1≤k≤n−1Y_{k} = S_{k} / S_{n}, 1 \leq k \leq n-1, have the same distribution as the order statistics of independent variables {Uk:1≤k≤n−1}\left\{ U_{k}: 1 \leq k \leq n-1\right\} which are uniformly distributed on (0,1)(0,1),

(b)

Zk=Xk/Sn,1≤k≤nZ_{k} = X_{k} / S_{n}, 1 \leq k \leq n, have the same joint distribution as the coordinates of a point (U1,…,Un)\left(U_{1}, \ldots , U_{n}\right) chosen uniformly at random on the simplex ∑r=1nur=1,ur≥0\sum_{r = 1}^{n} u_{r} = 1, u_{r} \geq 0 for all rr.

Conditional property of simple birth. Let X={X(t):t≥0}X = \left\{ X(t): t \geq 0\right\} be a simple birth process with rate λ\lambda, and let X(0)=1X(0) = 1. Let b≥1b \geq 1. Show that, conditional on the event {X(t)=b+1}\left\{ X(t) = b+1\right\}, the times of the bb births have the same distribution as the order statistics of a random sample of size bb from the density function

f(x)=λe−λ(t−x)1−e−λt,0≤x≤t f(x) = \frac{\lambda e^{-\lambda (t-x)}}{1-e^{-\lambda t}}, \quad 0 \leq x \leq t
?

Suppose that we are at time zero. Passengers arrive at a train station according to a Poisson process with intensity λ\lambda. Compute the expected value of the total waiting time of all passengers who have come to the station in order to catch a train that leaves at time tt.

?

Suppose that electrical pulses having i.i.d. random amplitudes A1,A2,…A_1, A_2, \ldots arrive at a counter in accordance with a Poisson process with intensity λ\lambda. The amplitude of a pulse is assumed to decrease exponentially, that is, if a pulse has amplitude AA upon its arrival, then its amplitude at time tt is Ae−αtAe^{-\alpha t}, where α\alpha is some positive parameter. We finally assume that the initial amplitudes of the pulses are independent of the Poisson process. Compute the expected value of the total amplitude at time tt.

?

Let {X(t),t≥0}\left\{ X(t), t \geq 0\right\} be a Poisson process and let {Tk,k≥1}\left\{ T_k, k \geq 1\right\} be the occurrence times. Suppose that we know that T3=1T_3=1 and that T1=xT_1=x, where 0<x<10 < x < 1. Our intuition then tells us that the conditional distribution of T2T_2 should be U(x,1)\mathrm{U}(x,1)-distributed. Prove that this is indeed the case, i.e., show that

T2∣T1=x,T3=1∈U(x,1)for0<x<1. T_2 \mid T_1=x, T_3=1 \in \mathrm{U}(x,1) \quad \text{for} \quad 0 < x < 1.
?

A train departs at time s>0s > 0. Starting from time t=0t=0, passengers arrive in a Poisson stream with intensity λ\lambda. Prove that the mean total waiting time of the passengers equals λs2/2\lambda s^{2}/2.

?

In a Poisson process one has observed four events in the interval (0,4](0,4]. Compute the probability that two occurred in (0,1](0,1] and the remaining two in (1,2](1,2].

?

Let {X(t)}\left\{ X(t)\right\} be a Poisson process with intensity λ\lambda and denote by T1T_{1} the time of the first event. Then, P(T1≤s)=1−exp⁡(−λs)\mathbb {P}\left(T_{1} \leq s\right) = 1-\exp (-\lambda s). Show that, for 0<s<t0 < s < t,

P(T1≤s∣X(t)=1)=s/t. \mathbb {P}\left(T_{1} \leq s \mid X(t)=1\right) = s/t .
?

Denote by vtv_{t} the random number of requests of a Poisson flow that arrive in the interval [0,t)[0, t). Prove that for t⩽T,k⩽nt \leqslant T, k \leqslant n,

P(vt=k∣vT=n)=Cnk(tT)k(1−tT)n−k. \mathbb {P}\left( v_{t} = k \mid v_{T} = n \right) = C_{n}^{k}\left(\frac{t}{T}\right)^{k}\left(1-\frac{t}{T}\right)^{n-k} .
?

Prove that if it is known that NN requests of a Poisson flow arrived on the segment [0,T],T>0[0, T], T > 0, then the flow of requests on this segment is a Bernoulli flow.

?

References

[ГюКу03] Гюнтер Н.М., Кузьмин Р.О. Сборник задач по высшей математике. 13-е изд., Издательство «Лань», 2003.
[Bas11] Bass R.F. Stochastic Processes. Cambridge University Press, 2011.
[Bil95] Billingsley P. Probability and Measure. 3rd ed., Wiley-Interscience, 1995.
[Cac89] Cacoullos T. Exercises in Probability. Springer-Verlag, 1989.
[Gal13] Gallager R.G. Stochastic Processes: Theory for Applications. Cambridge University Press, 2013.
[GrSt20] Grimmett G.R., Stirzaker D.R. Probability and Random Processes. 4th ed., Oxford University Press, 2020.
[Gut09] Gut A. An Intermediate Course in Probability. 2nd ed., Springer, 2009.
[КоФо03] Коршунов Д.А., Фосс С.Г. Сборник задач и упражнений по теории вероятностей. 2-е изд., Новосибирский государственный университет, 2003.
[LRS14] Lindgren G., Rootzén H., Sandsten M. Stationary Stochastic Processes for Scientists and Engineers. CRC Press (Taylor & Francis Group), 2014.
[ПУУ86] Прохоров А.В., Ушаков В.Г., Ушаков Н. Задачи по теории вероятностей. Наука. Главная редакция физико-математической литературы, 1986.